7. Sketch ray diagrams Ineatness counts] for a convex lens in which a. M 3-0.5 b. M 1.5
Answer #1

Imaze 5 Obiect 10 F

m = v/u = -0.5

v = -0.5u

where v = image distance and u = object distance.

Let’s assume u = 10 cm

then

v = -5 cm

(b)

10 Objet mage 15

m = v/u = -1.5

v = -1.5u

where v = image distance and u = object distance.

Let’s assume u = 10 cm

then

v = -15 cm